OpenStudy (calculusxy):

Medal Giveaway! After Richard gave 10% of his money to charity, he had $180 left. How much money did he originally have?

OpenStudy (loser66):

200

OpenStudy (calculusxy):

Explain?

OpenStudy (loser66):

let x is the original amount of money so, 10% x = 0.1x is the money he gave to charity therefore x - 0.1x =180 solve for x to get 200

OpenStudy (calculusxy):

How did I solve for to get 200? @Loser66

OpenStudy (loser66):

0.9x =180 x = 180/0.9 =200

OpenStudy (calculusxy):

Thanks